# SGD with Clipping is Secretly Estimating the Median Gradient

### v1

@@ Abstract
There are several applications of stochastic optimization where one can benefit from a robust estimate of the gradient. For example, domains such as distributed learning with corrupted nodes, the presence of large outliers in the training data, learning under privacy constraints, or even heavy-tailed noise due to the dynamics of the algorithm itself. Here we study SGD with robust gradient estimators based on estimating the median. We first consider computing the median gradient across samples, and show that the resulting method can converge even under heavy-tailed, state-dependent noise. We then derive iterative methods based on the stochastic proximal point method for computing the geometric median and generalizations thereof. Finally we propose an algorithm estimating the median gradient across iterations, and find that several well known methods – in particular different forms of clipping – are particular cases of this framework.
Machine Learning, ICML
[…]
figuret
@@ 1 Introduction
Many problems in machine learning can be represented by
@@
an optimization problem where $\bm{w}$ is a vector of parameters and $\ell:\mathbb{R}^{d}\to\mathbb{R}$ is a loss function. To tackle this problem, gradient-based optimization algorithms have been the de facto choice in many application domains (Nemirovsky & Yudin, 1983; Bottou, 2010; Bottou et al., 2018), where we are only given access to a noisy oracle of the gradient, i.e.,
$\bm{g}_{t}=\nabla\ell(\bm{w}_{t})+\bm{\xi}_{t}(\bm{w}_{t}).$
Here $t$ denotes the iterations and $\bm{\xi}_{t}(\bm{w}_{t})\in\mathbb{R}^{d}$ is noise vector that is sampled from a distribution that depends on the parameter $\bm{w}_{t}$. With this noise oracle, stochastic gradient descent (SGD) with learning rate $\eta_{t}$ is given by
@@
While SGD has proven useful in numerous applications, its performance heavily relies on the ‘quality’ of the noisy oracles $\bm{g}_{t}$, and can tragically degrade when the noise $\bm{\xi}_{t}$ has significant outliers, exhibits heavy tails (i.e., $\|\bm{\xi}_{t}\|$ can be very large with non-negligible probability), or has been corrupted in an adversarial way. We provide two examples.
@@
    Heavy-tailed gradients. Recent studies have provided theoretical and empirical evidence that heavy tails can naturally arise in stochastic optimization. From a theoretical perspective (Gurbuzbalaban et al., 2021; Hodgkinson & Mahoney, 2021; Pavasovic et al., 2023) showed that the parameter-dependent nature of $\bm{\xi}_{t}$ can result in heavy tails, even in very basic settings such as linear regression with Gaussian data. On the other hand, it has been empirically observed that the gradients for training transformer architectures on language tasks are more heavy-tailed compared to, for example, convolutional models for image data (Zhang et al., 2020b; Kunstner et al., 2023). While the presence of heavy-tails in gradients might be beneficial in certain settings (Simsekli et al., 2020), from an optimization perspective, it mostly introduces non-trivial challenges, which might even make the algorithm diverge unless additional care is taken (Zhang et al., 2020b; Gorbunov et al., 2020).
@@
    Corrupted nodes. A well-known problem in distributed learning is when some computation nodes are malicious and can communicate adversarial updates, which results in inaccurate gradient oracles that can significantly misguide the optimization procedure. Similar to the heavy-tailed setting, depending on how much malicious are the nodes, SGD can be impractical unless certain modifications are made (Mhamdi et al., 2018). We refer to (Karimireddy et al., 2021) and references therein for an overview of more robust aggregation techniques.
In order to make the optimization algorithms more robust to such noisy oracles, many techniques have been developed for different domains under various conditions. In the context of heavy-tailed gradients, clipping is often employed (Zhang et al., 2020a; Puchkin et al., 2023; Koloskova et al., 2023). In distributed learning robust aggregation rules are needed to protect against corrupted or malicious nodes. (Data et al., 2018; Karimireddy et al., 2021; Khirirat et al., 2023). This line of research is also closely related to error feedback in federated learning (Seide et al., 2014; Karimireddy et al., 2019; Richtárik et al., 2021) and learning under differential privacy constraints (Khirirat et al., 2023).
@@
Inspired by tools from robust statistics, some notions of median have been utilized in stochastic optimization as well (Acharya et al., 2022; Yin et al., 2018). One such approach is to compute a sample median over a sample of gradients. That is, let
$\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\xi}^{(i)}_{t}(\bm{w}_{t})\quad\mbox{for}\quad i=1,\dots,n,$
where $\{\bm{\xi}^{(i)}_{t}(\bm{w}_{t})\in\mathbb{R}^{d}\}_{i=1}^{n}$ is a set of i.i.d. noise vectors.
Given these sampled gradients, stochastic optimization algorithms typically aggregate them by computing their sample mean, or average $\bm{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\bm{g}_{t}^{(i)}$, and then use this $\bm{g}_{t}$ in (SGD). Alternatively, to improve robustness, one can aggregate these gradients $\{\bm{g}_{t}^{(i)}\}_{i=1}^{n}$ using a multidimensional variant of the median. While this approach has paved the way for powerful optimization algorithms in terms of robustness, it often introduces a significant computational burden, since a multivariate median needs to be computed at every iteration (Weiszfeld & Plastria, 2009; Vardi & Zhang, 2000).
@@
Motivation. Even though the clipping-based approaches developed in optimization theory and the median-based approaches inspired from robust statistics aim at solving similar tasks, they where designed in different mathematical frameworks under different conditions. In this study, our main goal is to bridge the gap between these two seemingly different branches and to bring a unified theoretical perspective that can shed further light on both directions by leveraging tools from both optimization and robust statistics.
Contributions. We first revisit median gradient-based optimization and consider a generic notion of median which include several multivariate medians that have been extensively used in practice. By exploiting the statistical properties of the sample median estimator, we show that SGD with (approximately computed) median gradients indeed yields a very powerful algorithm, which converges with $1/\sqrt{T}$ rate ($T$ being the number of iterations) even when the gradient oracles are heavy-tailed with diverging second-order moments. We further illustrate that, in addition to having infinite variance, the noise vectors $\bm{\xi}_{t}(\bm{w})$ can even have a multiplicative dependence on $\bm{w}$, a scenario which is highly challenging and cannot be directly covered by existing theoretical results, see e.g., (Wang et al., 2021; Zhang et al., 2020b; Puchkin et al., 2023).
Encouraged by this result, we then focus on the practical computation of the medians. To this end, we solve the optimization-based formulation of the median using Stochastic Proximal Point (SPP) (Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019; Toulis et al., 2020). As an alternative, we also consider iterative solvers based on the stochastic subgradient method.
We then depart from previous work by approximately computing the median gradient across iterations. That is, the median over $\{\bm{g}_{t}\}_{t=0}^{T}.$ We use this online median gradient within the SGD framework. When using the Stochastic Proximal Point method for computing the median across iterations, we recover several known clipping-based optimization algorithms as special cases. In particular, we recover different forms of clipping (Gorbunov et al., 2020; Zhang et al., 2020a), such as Clip21 for distributed learning (Khirirat et al., 2023), and signed gradient methods (Riedmiller & Braun, 1993; Bernstein et al., 2018). We also show that (non-clipped) heavy-ball momentum (Polyak, 1964) is in fact an online mean estimator.
@@
Our proposed framework allows to unify the different motivations for previously developed techniques, and illustrates that clipping-based optimization algorithms essentially runs a “hidden” estimation of the median across iterations. In particular, the observation that centered clipping (Karimireddy et al., 2021) and Clip21 are an iterative estimator of a geometric median appears to be new, and might have consequences for its application in distributed learning.
@@
We finally illustrate our theory on experiments where we compare the effectiveness of sample median, sample mean, and several online median estimators. Our results underline that for heavy-tailed noise, using the sample median is highly effective in contrast to the sample mean which is unstable and often does not converge. Our experiments also show that our online median estimates are robust, while being less expensive to compute. We further compare different clipping techniques for training transformer architectures on language modeling tasks, and show that they improve upon the performance of SGD with momentum.
[…]
@@ 2.1 Heavy-tailed Random Variables
A random variable $X$ is said to be heavy-tailed if the tails of its distribution decay slower than the tails of an exponential distribution. One important example, is the family of symmetric $\alpha$-stable distributions (Nolan, 2020, Def. 1.4), denoted $\mathcal{P}_{\alpha,\sigma}$: it is parameterized by a stability parameter $\alpha\in(0,2]$ and a scale parameter $\sigma>0$. For $\alpha=1$, it is the Cauchy distribution, while for $\alpha=2$ it is a Gaussian. The parameter $\sigma$ controls the spread of the distribution. When $\sigma=1$, we call the distribution standard and denote it by $\mathcal{P}_{\alpha}:=\mathcal{P}_{\alpha,1}$. The parameter $\alpha$ on the other hand controls the heaviness of the tails: for $\alpha<2$ and $X\sim\mathcal{P}_{\alpha}$ has infinite variance, whereas for $\alpha\leq 1$ even the mean of $X$ is infinite. However, for any $\alpha$, the median is equal to 00.
@@ 2.2 Mean Estimation and (Sample) Median
Estimating the mean of a distribution, given a finite number of samples, is one of the core problems studied in statistics. When the distribution in question has heavy-tails, or has significant outliers, the sample mean turns out to be a poor estimator. We refer to (Lugosi & Mendelson, 2019) for a survey of alternative techniques. In this section, we will introduce basic definitions for median.
In this section we denote by $z$ a real-value random value, and with $\bm{z}\in\mathbb{R}^{d}$ a random vector. We always denote the expectation with respect to the distribution of $z$ (or $\bm{z}$ respectively) by $\mathbb{E}$.
@@
In contrast to the mean, the median is always defined (though it may not be unique) (Cramér, 2016).
Since we are interested in approximating the median gradient, we need a notion of median that generalizes to random vectors $\bm{z}\in\mathbb{R}^{d}$. For this, we can define the $\ell_{p}$-median:
$\texttt{median}_{\ell_{p}}(\bm{z})\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}\left[\|\bm{m}-\bm{z}\|_{p}\right].$ (2)
This recovers several existing notions of a multivariate median: for $p=1$ we obtain a componentwise median (or $\ell_{1}$-median) and for $p=2$ we obtain the geometric median (or $\ell_{2}$-median) (Fréchet, 1948; Weiszfeld & Plastria, 2009; Minsker, 2015; Cohen et al., 2016). Clearly, for $d=1$ all notions conincide with the definition in (1).
@@
We remark that (3) is a special case of (2), by setting the distribution of $\bm{z}$ as a sum of Dirac measures $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{z}_{i}}$.
It is generally accepted that the sample median is a more robust estimator than the sample mean (Huber, 1981; Hampel et al., 2005). This is nicely illustrated by the notion of a breakdown point (Donoho & Huber, 1983; Lopuhaa & Rousseeuw, 1991), which is the smallest fraction of a sample that, if arbitrarily corrupted, can arbitrarily change the value of the estimator. The breakdown point of the sample median is roughly $\tfrac{1}{2}$, meaning that at least half of the samples need to be outliers for the median to diverge. The sample mean however can be arbitrarily corrupted by a single outlier, thus its breakdown point is one over $\tfrac{1}{n}$.
@@
More generally, for a function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ we will later consider the problem
@@
We assume that $\mathcal{D}$ is a closed, proper, convex function and that (4) admits a solution (see Appendix B for a discussion). For $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}$, the solution to (4) is clearly the mean $\mathbb{E}\left[\bm{z}\right]$, and for $\mathcal{D}=\|\cdot\|_{2}$, the solution is the geometric median as defined in (2). Problem (4) can also be seen as a special case of M-estimators (Huber, 1981, Sec. 3.2).

### v2

# Tracking the Median of Gradients with a Stochastic Proximal Point Method

@@ Abstract
There are several applications of stochastic optimization where one can benefit from a robust estimate of the gradient. For example, domains such as distributed learning with corrupted nodes, the presence of large outliers in the training data, learning under privacy constraints, or even heavy-tailed noise due to the dynamics of the algorithm itself. Here we study SGD with robust gradient estimators based on estimating the median.
We first derive iterative methods based on the stochastic proximal point method for computing the median gradient and generalizations thereof. Then we propose an algorithm estimating the median gradient across iterations, and find that several well known methods are particular cases of this framework. For instance, we observe that different forms of clipping allow to compute online estimators of the median of gradients, in contrast to (heavy-ball) momentum, which corresponds to an online estimator of the mean. Finally, we provide a theoretical framework for an algorithm computing the median gradient across samples, and show that the resulting method can converge even under heavy-tailed, state-dependent noise.
@@ 1 Introduction
Many problems in machine learning can be represented by an optimization problem
@@
where $\bm{w}$ is a vector of parameters and $\ell:\mathbb{R}^{d}\to\mathbb{R}$ is a loss function. To tackle this problem, gradient-based optimization algorithms have been the de-facto choice in many application domains (Nemirovsky & Yudin, 1983; Bottou, 2010; Bottou et al., 2018), where we are only given access to a noisy oracle of the gradient
$\bm{g}_{t}=\nabla\ell(\bm{w}_{t})+\bm{\xi}_{t}(\bm{w}_{t}).$ (1)
Here $t$ denotes the iterations and $\bm{\xi}_{t}(\bm{w}_{t})\in\mathbb{R}^{d}$ is a noise vector that is sampled from a distribution that depends on the parameter $\bm{w}_{t}$. With this noise oracle, stochastic gradient descent (SGD) with learning rate $\eta_{t}$ is given by
@@
While SGD has proven useful in numerous applications, its performance heavily relies on the ‘quality’ of the noisy oracles $\bm{g}_{t}$, and can tragically degrade when the noise $\bm{\xi}_{t}$ has significant outliers, exhibits heavy tails (that is, $\|\bm{\xi}_{t}\|$ is very large with non-negligible probability), or has been corrupted in an adversarial way. We provide two examples.
@@
    Heavy-tailed gradients. Recent studies have provided theoretical and empirical evidence that heavy tails can naturally arise in stochastic optimization. From a theoretical perspective Gurbuzbalaban et al. (2021); Hodgkinson & Mahoney (2021); Lehman et al. (2023) showed that the parameter-dependent nature of $\bm{\xi}_{t}$ can result in heavy tails, even in very basic settings such as linear regression with Gaussian data. On the other hand, it has been empirically observed that the gradients for training transformer architectures on language tasks are more heavy-tailed compared to, for example, convolutional models for image data (Zhang et al., 2020b; Kunstner et al., 2023). While the presence of heavy-tails in gradients might be beneficial in certain settings (Simsekli et al., 2020), from an optimization perspective, it mostly introduces non-trivial challenges, which might even make the algorithm diverge unless additional care is taken (Zhang et al., 2020b; Gorbunov et al., 2020).
@@
    Corrupted nodes. A well-known problem in distributed learning is when some computation nodes are malicious and can communicate adversarial updates, which results in inaccurate gradient oracles that can significantly misguide the optimization procedure. Similar to the heavy-tailed setting, depending on how malicious the nodes are, SGD can be impractical unless certain modifications are made (Mhamdi et al., 2018). We refer to Karimireddy et al. (2021) and references therein for an overview of more robust aggregation techniques.
In order to make the optimization algorithms more robust to such noisy oracles, many techniques have been developed for different domains under various conditions. In the context of heavy-tailed gradients, clipping is often employed (Zhang et al., 2020a; Puchkin et al., 2023; Koloskova et al., 2023). In distributed learning robust aggregation rules are needed to protect against corrupted or malicious nodes (Data et al., 2018; Karimireddy et al., 2021; Khirirat et al., 2023). This line of research is also closely related to error feedback in federated learning (Seide et al., 2014; Karimireddy et al., 2019; Richtarik et al., 2021) and learning under differential privacy constraints (Khirirat et al., 2023).
@@
Inspired by tools from robust statistics, some notions of median have been utilized in stochastic optimization as well (Yin et al., 2018; Alistarh et al., 2018; Acharya et al., 2022). One such approach is to compute the median over a finite sample of gradients. While this approach has paved the way for powerful optimization algorithms in terms of robustness, it often introduces a significant computational burden, since a multivariate median needs to be computed at every iteration (Weiszfeld & Plastria, 2009; Vardi & Zhang, 2000).
@@
Motivation. Clipping-based and error feedback approaches on the one hand, and median-based approaches on the other hand are motivated through different mathematical frameworks and appear to use different algorithmic tools. In this study, our main goal is to bridge the gap between these two seemingly different branches and to bring a unified theoretical perspective that can shed further light on both directions. The main tool to build this bridge between stochastic optimization and robust statistics will be the Stochastic Proximal Point method (SPP) (Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019; Toulis et al., 2020) for solving a later-specified gradient estimation problem.
Contributions. We first show how heavy-ball momentum (Polyak, 1964) can be seen as online estimator of the mean gradient using SPP. Here we use the term online in the sense that the method is given a different loss function at each step.
Encouraged by this result, we then focus on online estimation of the median gradient, using the same technique. Doing so, we recover several known clipping-based optimization algorithms from distributed learning as special cases, such as Clip21 (Khirirat et al., 2023), and a variant of centered clipping (Karimireddy et al., 2021). We also shown connections to sign-based gradient methods (Riedmiller & Braun, 1993).
@@
Our proposed framework allows to unify the different motivations for previously developed techniques, and illustrates that clipping-based optimization algorithms essentially run a loose estimation of the median across iterations. In particular, the observation that centered clipping and Clip21 are iterative estimators of a geometric median appears to be new, and might have consequences for its application in distributed learning.
We theoretically analyze these online gradient estimation methods in a simplified setting: when allowing multiple gradient samples per iteration, we show that SGD with (approximately computed) median gradients indeed yields a very powerful algorithm, which converges with $1/\sqrt{T}$ rate ($T$ being the number of iterations) even when the gradient oracles are heavy-tailed with diverging second-order moments. We further illustrate that, in addition to having infinite variance, the noise vectors $\bm{\xi}_{t}(\bm{w})$ can even have a multiplicative dependence on $\bm{w}$, a scenario which is highly challenging and cannot be directly covered by existing theoretical results, see e.g., Wang et al. (2021); Zhang et al. (2020b); Puchkin et al. (2023).
@@
We finally illustrate our theory on synthetic least-squares experiments where we compare the effectiveness of sample median, sample mean, and several online median estimators. Our results underline that for heavy-tailed noise, using the sample median is highly effective in contrast to the sample mean which is unstable and often does not converge. Our experiments also show that our online median estimates are robust, and require only a single sample per iteration, making it a less expensive alternative to the sample median. We further compare different clipping techniques for training transformer architectures on language modeling tasks, and show that they can improve upon the performance of SGD with momentum, however the gap is relatively small.
Notation. Throughout the paper, $\|\cdot\|_{p}$ denotes the $\ell_{p}$-norm, given by $\|\bm{z}\|_{p}:=\left(\sum_{i=1}^{d}|\bm{z}|_{i}^{p}\right)^{\frac{1}{p}}$ when $p\in[1,\infty)$, and $\|\bm{z}\|_{\infty}=\max\limits_{i=1,\dots,d}|\bm{z}_{i}|$ when $p=\infty$. When $p=2$ we recover the usual Euclidean $\ell_{2}$-norm $\|\cdot\|_{2}$, which we will often simply denote by $\|\cdot\|$. For a matrix $\bm{A}$, we denote its Frobenius norm by $\|\bm{A}\|_{F}:=\sqrt{\sum_{i,j}\bm{A}_{ij}^{2}}$. For $n\in\mathbb{N}_{+}$, we denote the set $n]:=\{1,\dots,n\}$.
@@ 2.1 Heavy-tailed Random Variables
A random variable $X$ is said to be heavy-tailed if the tails of its distribution decay slower than the tails of an exponential distribution. One important example is the family of symmetric $\alpha$-stable distributions $\mathcal{P}_{\alpha,\sigma}$ (Nolan, 2020, Def. 1.4), parameterized by a stability index $\alpha\in(0,2]$ and a scale $\sigma>0$. When $\alpha=1$, this is the Cauchy distribution; when $\alpha=2$, it reduces to a Gaussian. The parameter $\sigma$ controls the spread of the distribution. When $\sigma=1$, we call the distribution standard and denote it by $\mathcal{P}_{\alpha}:=\mathcal{P}_{\alpha,1}$. The parameter $\alpha$ on the other hand controls the heaviness of the tails: for $\alpha<2$ the random variable $X\sim\mathcal{P}_{\alpha}$ has infinite variance, whereas for $\alpha\leq 1$ even the mean of $X$ is infinite. However, for any $\alpha$, the median is equal to 0.
@@ 2.2 Mean Estimation and (Sample) Median
Estimating the mean of a distribution, when given a finite number of samples, is one of the core problems studied in statistics. If the distribution in question has heavy-tails, or has significant outliers, the sample mean turns out to be a poor estimator. Lugosi & Mendelson (2019) give a survey of alternative and more robust techniques: one such alternative is the median, for which we will introduce basic definitions next.
Here, we denote by $z$ a real-value random value, and the boldface $\bm{z}\in\mathbb{R}^{d}$ a random vector. We always denote the expectation with respect to the distribution of $z$ (or $\bm{z}$ respectively) by $\mathbb{E}$.
@@
In contrast to the mean, the median is always defined (though it may not be unique) (Cramér, 2016). Since we are interested in approximating the median gradient, we need a notion of median that generalizes to random vectors $\bm{z}\in\mathbb{R}^{d}$. For this, we can define the $\ell_{p}$-median:
$\texttt{median}_{\ell_{p}}(\bm{z})\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}\left[\|\bm{m}-\bm{z}\|_{p}\right],$ (3)
where $\|\bm{m}\|_{p}=\left(\sum_{j=1}^{d}\bm{m}_{j}^{p}\right)^{\frac{1}{p}}$ is the $\ell_{p}$ norm. This recovers several existing notions of a multivariate median: for $p=1$ we obtain a componentwise median (or $\ell_{1}$-median) and for $p=2$ we obtain the geometric median (or $\ell_{2}$-median) (Fréchet, 1948; Weiszfeld & Plastria, 2009; Minsker, 2015; Cohen et al., 2016). For $d=1$ and every $p\in[1,\infty)$, the $\ell_{p}$-median coincides with the one-dimensional median in (2).
@@
Clearly (4) is a special case of (3), by setting the distribution of $\bm{z}$ as the empirical measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{z}_{i}}$.
It is generally accepted that the sample median is a more robust estimator than the sample mean (Huber, 1981; Hampel et al., 2005). This is nicely illustrated by the notion of a breakdown point (Donoho & Huber, 1983; Lopuhaa & Rousseeuw, 1991), which is the smallest fraction of a sample that, if arbitrarily corrupted, can arbitrarily change the value of the estimator. The breakdown point of the sample median is roughly $\tfrac{1}{2}$, meaning that at least half of the samples need to be outliers for the median to diverge. The sample mean however can be arbitrarily corrupted by a single outlier, thus its breakdown point is $\tfrac{1}{n}$.
@@
More generally, for a function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ we will consider the problem
@@
We assume that $\mathcal{D}$ is a closed, proper, convex function and that (5) admits a solution (see Appendix˜B for a discussion). For $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}$, the solution to (5) is the mean $\mathbb{E}\left[\bm{z}\right]$ (if it exists), and for $\mathcal{D}=\|\cdot\|_{2}$ the solution is the geometric median as defined in (3). Problem (5) can be seen as a special case of M-estimators (Huber, 1981, Sec. 3.2).